### v1

# A refinement of the Pontryagin-Thom theorem for unstable Thom spectra and its applications

Abstract.
The Pontryagin-Thom construction provides a fundamental link between cobordism groups and the homotopy groups of Thom spectra. Our main result refines this theorem, providing a more explicit geometric interpretation of the homotopy groups of unstable Thom spectra. Building on this result, we show that previously unknown cobordism groups can be expressed as homotopy groups of unstable Thom spectra. Furthermore, using the Smith homomorphism, we compute these groups. As applications, we determine the values of $n$ for which there exists a Spin manifold with boundary $S^{n}$ admitting a line subbundle orthogonal to the boundary, and provide a precise characterization of the cobordism group introduced by Bais, May Custodio, and Torres.
Key words and phrases:
cobordism group, spectrum
2020 Mathematics Subject Classification:
57R90, 55P42
1. Introduction
In topology, the classification of manifolds is a fundamental and important problem. The cobordism group, introduced by Thom, is one of the classical objects of this study. By the Pontryagin-Thom construction, this cobordism group is known to be isomorphic to the homotopy groups of the corresponding Thom spectrum [12].
An analogue of Thom’s cobordism group was introduced by Reinhart, defined by a stronger equivalence relation [9]. Reinhart determined this group by using a natural surjection onto the Thom’s original cobordism group. Later, Ebert [6, Appendix A] noted that this group is also isomorphic to the homotopy groups of a corresponding unstable Thom spectrum and, by using a long exact sequence of homotopy groups, obtained the same result as Reinhart.
The theorem used by Ebert [6] is the following, which summarizes [10, Chapter IV, §7]. Before stating the theorem, we define the relevant cobordism group.
Henceforth, we denote the Thom spectrum of a virtual vector bundle $V\to X$ by $X^{V}$.
Definition 1.1.
Let $V\to X$ be a virtual vector bundle of rank $n\in{\mathbb{Z}}_{\geq 0}$. We define the monoid denoted by $\Omega(X;V)$, as the set of equivalence classes of triples $(M,g,\Phi)$ consisting of:
  • a closed $n$-manifold $M$,
  • a continuous map $g:M\to X$, and
  • a stable vector bundle isomorphism $\Phi:TM\cong g^{*}V$ (i.e., there exists a sufficiently large positive integer $N$ and a bundle isomorphism $\Phi:TM\oplus\mathbb{R}^{N}\cong g^{V}\oplus\mathbb{R}^{N}$).
The equivalence relation is defined as follows. A direct cobordism between two triples $(M_{0},g_{0},\Phi_{0})$ and $(M_{1},g_{1},\Phi_{1})$ is a triple $(N,h,\Psi)$ consisting of:
  • a compact $(n+1)$-manifold NN which is a bordism from $M_{0}$ to $M_{1}$,
  • a continuous map $h:N\to X$ such that $h|_{M_{i}}=g_{i}$, and
  • a stable bundle isomorphism $\Psi:TN\cong h^{*}V\oplus{\mathbb{R}}$ whose restriction to $M_{i}$ coincides with the isomorphism
    $TN|_{M_{i}}\cong TM_{i}\oplus{\mathbb{R}}\stackrel{{\scriptstyle\Phi_{i}}}{{\cong}}g_{i}^{*}V\oplus{\mathbb{R}}.$
Here, the isomorphism $TM_{i}\oplus\mathbb{R}\cong TN$ identifies the trivial bundle with the outward normal vector for $i=0$ and the inward normal vector for $i=1$.
Since this direct cobordism relation is reflexive and transitive but not necessarily symmetric, we define the equivalence relation on these triples to be the one generated by the direct cobordism relation. That is, two triples are equivalent if they can be connected by a finite sequence of direct cobordisms forming a “zigzag”.
Here, the addition in the monoid is given by the disjoint union of manifolds.
Theorem 1.2.
Let $V\to X$ be a virtual vector bundle of rank $n\in{\mathbb{Z}}_{\geq 0}$. Then, there is a natural isomorphism
$\Omega(X;V)\cong\pi_{0}(X^{-V}).$
The correspondence is as follows. Given a triple $M,g,\Phi]$, the corresponding element in $\pi_{0}(X^{-V})$ is the composition
$\Sigma^{\infty}\mathbb{S}^{0}\xrightarrow{\operatorname{PT}_{M}}M^{-TM}\xrightarrow{g,\Phi}X^{-V}$
The converse correspondence is given by Thom’s transversality theorem.
This theorem is highly significant in that it translates the Thom spectrum into the language of manifolds. However, it has the disadvantage that the “stable vector bundle isomorphism” appearing in the equivalence relation is difficult to handle. We refine this theorem and express the equivalence relation in a more explicit form.
The relevant cobordism group for our first main result is defined as follows.
Definition 1.3.
Let $m$ be a positive integer, $n$ be a non-negative integer, and let $\xi\to X$ be a vector bundle of rank $n+m$. We define the monoid denoted by $\Omega^{\prime}(X;\xi,m)$, as the set of equivalence classes of triples $(M,g,\Phi)$ consisting of:
  • a closed smooth $n$-manifold $M$,
  • a continuous map $g:M\to X$, and
  • a bundle isomorphism $\phi:TM\oplus{\mathbb{R}}^{m}\cong g^{*}\xi$.
The equivalence relation is defined as follows. A direct cobordism between two triples $(M_{0},g_{0},\phi_{0})$ and $(M_{1},g_{1},\phi_{1})$ is a triple $(N,h,\psi)$ consisting of:
  • a compact $(n+1)$-manifold NN which is a bordism from $M_{0}$ to $M_{1}$,
  • a continuous map $h:N\to X$ such that $h|_{M_{i}}=g_{i}$, and
  • a bundle isomorphism $\psi:TN\oplus{\mathbb{R}}^{m-1}\cong h^{*}\xi$ whose restriction to $M_{i}$ coincides with the isomorphism
    $TN\oplus{\mathbb{R}}^{m-1}|_{M_{i}}\cong TM_{i}\oplus{\mathbb{R}}^{m}\stackrel{{\scriptstyle\phi_{i}}}{{\cong}}g_{i}^{*}\xi.$
Here, the isomorphism $TM_{i}\oplus\mathbb{R}\cong TN$ identifies the trivial bundle with the outward normal vector for $i=0$ and the inward normal vector for $i=1$.
Since this direct cobordism relation is reflexive and transitive but not necessarily symmetric, we define the equivalence relation on these triples to be the one generated by the direct cobordism relation. That is, two triples are equivalent if they can be connected by a finite sequence of direct cobordisms forming a “zigzag”.
Here, the addition in the monoid is given by the disjoint union of manifolds.
Theorem A.
Let $m$ be a positive integer, $n$ be a non-negative integer, and let $\xi\to X$ be a vector bundle of rank $n+m$. Then, the natural monoid homomorphism
$\lambda:\Omega^{\prime}(X;\xi,m)\to\Omega(X;\xi-\mathbb{R}^{m})$
defined by $M,g,\phi]\mapsto[M,g,\phi]$ induces an isomorphism
$\Omega^{\prime}(X;\xi,m)\xrightarrow{\cong}\pi_{-m}(X^{-\xi}).$
Moreover, the direct cobordism relation defined in Definition 1.3 is an equivalence relation except when $n=1$.
Furthermore, Ebert [6] succeeded in obtaining information about the homotopy groups using a fibration of spectra, in which the homotopy groups of the corresponding Thom spectrum appear. On the other hand, the Smith homomorphism is also known as a method for constructing fibrations of Thom spectra [5]. In this paper, we use this Smith homomorphism to determine previously unknown cobordism groups. We begin by presenting three $3\times 3$ commutative diagrams involving the Smith homomorphism used in this study.
Theorem B1.
Let $n$ be a positive integer, $L^{O}_{n}$ be the canonical rank $n$ vector bundle over $BO(n)$, and $L^{SO}_{n}$ be the canonical rank $n$ oriented vector bundle over $BSO(n)$. Then, we obtain the commutative diagram
$BSO(n)^{-(L^{SO}_{n}\oplus{\mathbb{R}}^{2})}$$BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n}\oplus{\mathbb{R}})}$$BO(n)^{-(L^{O}_{n}\oplus{\mathbb{R}})}$$BSO(n+1)^{-(L^{SO}_{n+1}\oplus{\mathbb{R}})}$$BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})}$$BO(n+1)^{-(L^{O}_{n+1}\oplus{\mathbb{R}})}$$BSO(n+1)^{-{\mathbb{R}}}$$BO(n+1)^{-\mathrm{det}L^{O}_{n+1}}$$BO(n+1)^{0}$}
where every row and every column is a Smith homomorphism of certain Thom spectra.
Theorem B2.
Let $n$ be a positive integer, $L^{\mathrm{Spin}}_{n}$ be the canonical rank $n$ spin vector bundle over $B\mathrm{Spin}(n)$, and $L^{\mathrm{Pin}^{-}}_{n}$ be the canonical rank $n$ $\mathrm{Pin}^{-}$ vector bundle over $B\mathrm{Pin}^{-}(n)$. Then, we obtain the commutative diagram
$B\mathrm{Spin}(n)^{-(L^{\mathrm{Spin}}_{n}\oplus{\mathbb{R}}^{2})}$$B\mathrm{Pin}^{-}(n)^{-(L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n}\oplus{\mathbb{R}})}$$B\mathrm{Pin}^{-}(n)^{-(L^{\mathrm{Pin}^{-}}_{n}\oplus{\mathbb{R}})}$$B\mathrm{Spin}(n+1)^{-(L^{\mathrm{Spin}}_{n+1}\oplus{\mathbb{R}})}$$B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1})}$$B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus{\mathbb{R}})}$$B\mathrm{Spin}(n+1)^{-{\mathbb{R}}}$$B\mathrm{Pin}^{-}(n+1)^{-\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1}}$$B\mathrm{Pin}^{-}(n+1)^{0}$}
where every row and every column is a Smith homomorphism of certain Thom spectra.
Theorem B3.
Let $n$ be a positive integer, $L^{\mathrm{Spin}}_{n}$ be the canonical rank $n$ Spin vector bundle over $B\mathrm{Spin}(n)$, $L^{\mathrm{Spin}^{c}}_{n}$ be the canonical rank $n$ $\mathrm{Spin}^{c}$ vector bundle over $B\mathrm{Spin}^{c}(n)$, and $\mathcal{O}_{n}$ be the pullback of the canonical complex line bundle via the natural map $B\mathrm{Spin}^{c}(n)\to BSO(n)\times BU(1)\xrightarrow{pr_{2}}BU(1)$. Then, we obtain the commutative diagram
$B\mathrm{Spin}(n)^{-(L^{\mathrm{Spin}}_{n}\oplus{\mathbb{R}}^{3})}$$B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n}\oplus{\mathbb{R}})}$$B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus{\mathbb{R}})}$$B\mathrm{Spin}(n+1)^{-(L^{\mathrm{Spin}}_{n+1}\oplus{\mathbb{R}}^{2})}$$B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})}$$B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus{\mathbb{R}})}$$B\mathrm{Spin}(n+1)^{-{\mathbb{R}}^{2}}$$B\mathrm{Spin}^{c}(n+1)^{-\mathcal{O}_{n+1}}$$B\mathrm{Spin}^{c}(n+1)^{0}$}
where every row and every column is a Smith homomorphism of certain Thom spectra.
Next, we define three cobordism groups that have not been previously explored and use the mentioned three diagrams of the Smith homomorphism to extract information regarding their group structures. The three new groups addressed in this study are defined as follows.
Definition 1.4.
Let $n$ be a non-negative integer. We define the monoid $\Omega^{\text{SO, line}}_{n}$ as the set of equivalence classes of pairs $(M,\eta)$ consisting of:
  • a closed oriented $n$-manifold $M$, and
  • a line subbundle $\eta\subset TM\oplus{\mathbb{R}}$.
The equivalence relation is defined as follows. A direct cobordism between two pairs $(M_{0},\eta_{0})$ and $(M_{1},\eta_{1})$ is a pair $(N,\zeta)$ consisting of:
  • a compact oriented $(n+1)$-manifold NN which is a bordism from $M_{0}$ to $M_{1}$, and
  • a line subbundle $\zeta\subset TN$ whose restriction to $M_{i}$ coincides with $\eta_{i}$ under the identification $TN|_{M_{i}}\cong TM_{i}\oplus{\mathbb{R}}$.
Here, the isomorphism $TM_{i}\oplus{\mathbb{R}}\cong TN|_{M_{i}}$ identifies the trivial bundle with the outward normal vector for $i=0$ and the inward normal vector for $i=1$.
This direct cobordism relation is an equivalence relation except when $n=1$. In the case $n=1$, we define two elements to be equivalent if they can be connected by a finite sequence of direct cobordisms, forming a “zigzag”.
Here, the addition in the monoid is given by the disjoint union of manifolds.
Definition 1.5.
Let $n$ be a non-negative integer. We define the monoid $\Omega^{\text{Spin, line}}_{n}$ as the set of equivalence classes of pairs $(M,\eta)$ consisting of:
  • a closed Spin $n$-manifold $M$, and
  • a line subbundle $\eta\subset TM\oplus{\mathbb{R}}$.
The equivalence relation is defined as follows. A direct cobordism between two pairs $(M_{0},\eta_{0})$ and $(M_{1},\eta_{1})$ is a pair $(N,\zeta)$ consisting of:
  • a compact Spin $(n+1)$-manifold NN which is a bordism from $M_{0}$ to $M_{1}$, and
  • a line subbundle $\zeta\subset TN$ whose restriction to $M_{i}$ coincides with $\eta_{i}$ under the identification $TN|_{M_{i}}\cong TM_{i}\oplus{\mathbb{R}}$.
Here, the isomorphism $TM_{i}\oplus{\mathbb{R}}\cong TN|_{M_{i}}$ identifies the trivial bundle with the outward normal vector for $i=0$ and the inward normal vector for $i=1$.
This direct cobordism relation is an equivalence relation except when $n=1$. In the case $n=1$, we define two elements to be equivalent if they can be connected by a finite sequence of direct cobordisms, forming a “zigzag”.
Here, the addition in the monoid is given by the disjoint union of manifolds.
Definition 1.6.
Let $n$ be a non-negative integer. We define the monoid $\Omega^{\text{Spin, plane}}_{n}$ as the set of equivalence classes of pairs $(M,\eta)$ consisting of:
  • a closed Spin $n$-manifold $M$, and
  • a oriented rank 2 subbundle $\eta\subset TM\oplus{\mathbb{R}}^{2}$.
The equivalence relation is defined as follows. A direct cobordism between two pairs $(M_{0},\eta_{0})$ and $(M_{1},\eta_{1})$ is a pair $(N,\zeta)$ consisting of:
  • a compact Spin $(n+1)$-manifold NN which is a bordism from $M_{0}$ to $M_{1}$, and
  • a oriented rank 2 subbundle $\zeta\subset TN\oplus{\mathbb{R}}$ whose restriction to $M_{i}$ coincides with $\eta_{i}$ under the identification $TN|_{M_{i}}\oplus{\mathbb{R}}\cong TM_{i}\oplus{\mathbb{R}}^{2}$.
Here, the isomorphism $TM_{i}\oplus{\mathbb{R}}\cong TN|_{M_{i}}$ identifies the trivial bundle with the outward normal vector for $i=0$ and the inward normal vector for $i=1$.
This direct cobordism relation is an equivalence relation except when $n=1$. In the case $n=1$, we define two elements to be equivalent if they can be connected by a finite sequence of direct cobordisms, forming a “zigzag”.
Here, the addition in the monoid is given by the disjoint union of manifolds.
Theorem C1.
Let $n$ be a non-negative integer. There is a natural isomorphism
$\Omega^{\textnormal{SO, line}}_{n}\cong\Omega^{\prime}(BO(n);L^{O}_{n}\oplus\mathrm{det}L^{O}_{n},1)\cong\pi_{-1}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})}).$
Furthermore, there exists a natural surjective group homomorphism
$\epsilon:\Omega^{\textnormal{SO, line}}_{n}\to\Omega^{O}_{n-1}\oplus\Omega^{SO}_{n}$
defined by $\epsilon([M,\eta])=([s^{-1}(0)],[M])$, where $s$ is a generic section of $\eta$ transverse to the zero-section. The kernel of $\epsilon$ is given by
$\textrm{Ker}\,\epsilon\cong\begin{cases}{\mathbb{Z}}/2{\mathbb{Z}},&\text{if }n\equiv 1\pmod{4}\\ 0,&\text{if }n\equiv 0,2,3\pmod{4}\end{cases}$.
Theorem C2.
Let $n$ be a non-negative integer. There is a natural isomorphism
$\Omega^{\textnormal{Spin, line}}_{n}\cong\Omega^{\prime}(B\mathrm{Pin}^{-}(n);L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n},1)\cong\pi_{-1}(B\mathrm{Pin}^{-}(n)^{-(L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n})}).$
Furthermore, there exists a natural surjective group homomorphism
$\epsilon:\Omega^{\textnormal{Spin, line}}_{n}\to\Omega^{O}_{n-1}\oplus\Omega^{SO}_{n}$
defined by $\epsilon([M,\eta])=([s^{-1}(0)],[M])$, where $s$ is a generic section of $\eta$ transverse to the zero-section. The kernel of $\epsilon$ is given by
$\textrm{Ker}\,\epsilon\cong\begin{cases}{\mathbb{Z}}/2{\mathbb{Z}},&\text{if }n\equiv 1,3,4,5\pmod{8}\\ 0,&\text{if }n\equiv 0,2,6,7\pmod{8}\end{cases}$.
Theorem C3.
Let $n$ be a non-negative integer. There is a natural isomorphism
$\Omega^{\textnormal{Spin, plane}}_{n}\cong\Omega^{\prime}(B\textnormal{Spin}^{c}(n);L^{\textnormal{Spin}^{c}}_{n}\oplus\mathcal{O}_{n},2)\cong\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})}).$
Furthermore, there exists a natural surjective group homomorphism
$\epsilon:\Omega^{\textnormal{Spin, plane}}_{n}\to\Omega^{\textnormal{Spin}^{c}}_{n-2}\oplus\Omega^{\textnormal{Spin}}_{n}$
defined by $\epsilon([M,\eta])=([s^{-1}(0)],[M])$, where $s$ is a generic section of $\eta$ transverse to the zero-section. The kernel of $\epsilon$ is given by
[…]
As corollaries of our main theorems, we present two applications. The first application determines the non-negative integers $n$ for which there exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact Spin manifold $M$ with boundary $S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta$ is orthogonal to the boundary on $S^{n}$. The second application provides an alternative proof of the main result in [1]. In [1, Section 6], Question 2 asks for a description of the cobordism group defined in [1] in terms of the homotopy groups of a spectrum. Our work offers a solution to this question. Furthermore, using this description, we resolve the case $n\equiv 1\pmod{4}$, which was left open in [1].
Acknowledgements
The author thanks Takuya Sakasai for his support and helpful comments. This research was supported by FMSP, WINGS Program, the University of Tokyo.
2. Thom spectra and Pontryagin-Thom Construction
In this section, we review the Pontryagin-Thom construction, which provides a fundamental link between geometry and homotopy theory.
Begin by recalling, we recall the classical Pontryagin-Thom construction [12]. Let $M^{k}$ be a closed oriented $k$-dimensional smooth manifold embedded in $\mathbb{R}^{k+n}$ (or $S^{k+n}$). Let $\nu(M)$ be the normal bundle of this embedding. The orientation is naturally induced on this bundle. The Thom space $M^{\nu(M)}$ of the bundle $\nu(M)$ is defined as the quotient space $D(\nu)/S(\nu)$, where $D(\nu)$ and $S(\nu)$ are the disk bundle and sphere bundle of $\nu$, respectively.
The Pontryagin-Thom construction gives a “collapse map”
$f\colon S^{k+n}\to M^{\nu(M)}.$
This map is defined by collapsing the complement of a tubular neighborhood of $M$ in $S^{k+n}$ to the base point of the Thom space.
The construction leads to the main theorem, which relates oriented cobordisms to the stable homotopy groups of the Thom spectra of canonical oriented vector bundles.
Theorem 2.1 (Pontryagin-Thom Isomorphism).
Let $\Omega_{k}^{\mathrm{SO}}$ be the $k$-dimensional oriented cobordism group. There is an isomorphism
$\Phi\colon\Omega_{k}^{\mathrm{SO}}\xrightarrow{\cong}\pi_{k+n}(BSO(n)^{L_{n}^{\mathrm{SO}}})\quad(\text{for }n\text{ large enough})$
where $\pi_{k+n}(BSO(n)^{L_{n}^{\mathrm{SO}}})$ is the $(k+n)$-th homotopy group of the Thom space of the canonical rank $n$ oriented bundle $L_{n}^{\mathrm{SO}}$.
These framework can be extended to the setting of Thom spectra, which generalize Thom spaces. For definitions of the Thom spectrum associated with a virtual vector bundle and the generalized Pontryagin-Thom construction for Thom spectra, we refer the reader to [5, Section2] and [10].
3. The proof of theorem A
In this section, we prove Theorem A. The proof relies on obstruction theory, as also mentioned in [6, Remark A.3].
We first prove the following two lemmas, which are purely regarding vector bundles.
Lemma 3.1.
Let $M$ be an $n$-dimensional finite CW complex, and let $\xi_{1}$ and $\xi_{2}$ be $m$-dimensional vector bundles over $M$. If $m\geq n+1$ and $\xi_{1}\oplus{\mathbb{R}}$ and $\xi_{2}\oplus{\mathbb{R}}$ are isomorphic as vector bundles over $M$, then $\xi_{1}$ and $\xi_{2}$ are also isomorphic.
Proof.
We begin by fixing an isomorphism $\Phi:\xi_{1}\oplus{\mathbb{R}}\to\xi_{2}\oplus{\mathbb{R}}$. Our goal is to deform $\Phi$ over each cell of $M$ until it takes the form of an isomorphism between $\xi_{1}$ and $\xi_{2}$, followed by the identity $1\in GL(1)$ on the trivial line factor, i.e., $\phi\oplus 1\in GL(m+1)$.
This deformation process corresponds to extending a lift of the natural map $GL(m)\to GL(m+1)$ defined on the boundary $\partial D^{r}$ to the entire cell $D^{r}$. The obstruction to lifting lies in the relative homotopy group $\pi_{r}(GL(m+1),GL(m))\cong\pi_{r}(S^{m})$. Since $\pi_{r}(S^{m})=0$ for $r\leq m-1$, and our assumption $m\geq n+1$ implies $n\leq m-1$ when considering the $n$-dimensional skeleton of $M$, the obstruction vanishes for all cells of dimension $r\leq n$.
Therefore, the deformation can be performed over the entire space $M$, resulting in an isomorphism of the form $\phi\oplus 1$. This implies that $\xi_{1}$ and $\xi_{2}$ are isomorphic as vector bundles over $M$. ∎
Lemma 3.2.
Let $M$ be an $n$-dimensional finite CW complex, and let $\xi$ be an $m$-dimensional vector bundle over $M$. There is a natural bundle map from the $GL(m)$-bundle $\mathrm{Isom}(\xi)$ over $M$ to the $GL(m+1)$-bundle $\mathrm{Isom}(\xi\oplus{\mathbb{R}})$ over $M$, which is obtained by taking the direct sum with $1\in GL(1)$. Let $\Gamma(\mathrm{Isom}(\xi))$ denote the set of homotopy classes of sections of $\mathrm{Isom}(\xi)$.
Then, the map $\Gamma(\mathrm{Isom}(\xi))\to\Gamma(\mathrm{Isom}(\xi\oplus{\mathbb{R}}))$ induced by the bundle map is surjective when $m=n+1$, and is bijective when $m\geq n+2$.
Proof.
Surjectivity follows from the same argument as in the proof of Lemma 3.1. We now consider the case where $m\geq n+2$ and prove injectivity.
Let $\phi_{0}$ and $\phi_{1}$ be automorphisms of $\xi$. Assume that $\phi_{0}\oplus 1$ and $\phi_{1}\oplus 1$ are homotopic as automorphisms of $\xi\oplus{\mathbb{R}}$. This assumption means there exists an automorphism $\Phi$ of the pullback bundle $\mathrm{pr}_{1}^{\ast}\xi\oplus{\mathbb{R}}$ over $M\times[0,1]$ such that the restriction of $\Phi$ to $M\times\{i\}$ is $\phi_{i}\oplus 1$ for $i=0,1$.
We can now deform this automorphism $\Phi$ on the interior $M\times(0,1)$ following the strategy used in the proof of Lemma 3.1. Since $M\times[0,1]$ is an $(n+1)$-dimensional CW complex, and the dimension of the bundle is $m$, the condition for the vanishing of the obstruction to the deformation is that the dimension of the base space is less than or equal to m1m-1. Since $m\geq n+2$, we have $m-1\geq n+1$. Thus, the obstruction to performing the deformation over the $(n+1)$-dimensional space $M\times[0,1]$ vanishes.
Therefore, $\Phi$ can be deformed so that its restriction to $M\times\{t\}$ is of the form $\phi_{t}\oplus 1$ for all $t\in[0,1]$, where $\phi_{t}$ is an automorphism of $\xi$. This shows that $\phi_{0}$ and $\phi_{1}$ are homotopic as automorphisms of $\xi$. This establishes injectivity. ∎
Proof of Theorem A.
We begin by proving that $\lambda$ is an isomorphism. Surjectivity of $\lambda$ follows from Lemma 3.2. More precisely, let $\Phi$ be a stable isomorphism between $TM\oplus{\mathbb{R}}^{m}$ and $g^{*}\xi$, which means $\Phi$ is an isomorphism between $TM\oplus{\mathbb{R}}^{m+N}$ and $g^{*}\xi\oplus{\mathbb{R}}^{N}$ for some sufficiently large integer $N\geq 0$. Since the rank of $TM\oplus{\mathbb{R}}^{m}$ is $n+m\geq n+1$, and by the stability hypothesis of Lemma 3.2, there exists an isomorphism $\phi$ between $TM\oplus{\mathbb{R}}^{m}$ and $g^{*}\xi$ such that $\phi\oplus 1_{{\mathbb{R}}^{N}}$ is homotopic to $\Phi$. If we take this $\phi$ as the bundle isomorphism, we have $\lambda([M,g,\phi])=[M,g,\Phi]$. This proves surjectivity.
Since $\lambda$ is a surjective homomorphism from a commutative monoid to an abelian group, to prove injectivity, it suffices to show that if $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, then $M,g,\phi]=0\in\Omega^{\prime}(X;\xi,m)$. This follows from the logic that if $\lambda(a)=\lambda(b)$, we can choose an element $c$ such that $\lambda(c)=-\lambda(a)$. Then, $\lambda(a+c)=\lambda(b+c)=0$. If the implication holds (i.e., mapping to zero implies being zero), we have $a+c=b+c=0$, which implies $a=a+(b+c)=(a+c)+b=0+b=b$.
Since $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, there exist a sufficiently large positive integer $N$, a sequence of $n$-dimensional closed manifolds $M=M_{0},M_{1},\ldots,M_{l}$ with $M_{l+1}=\emptyset$, continuous maps $g_{i}:M_{i}\to X$, and bundle isomorphisms $\Phi_{i}$ between $TM_{i}\oplus{\mathbb{R}}^{m+N}$ and $g_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$. Note that we set $g_{0}=g$ and $\Phi_{0}=\phi\oplus{\mathbb{R}}^{N}$.
Furthermore, there exist a sequence of $(n+1)$-dimensional compact manifolds $N_{0},N_{1},\ldots,N_{l}$, maps $h_{i}:N_{i}\to X$, and isomorphisms $\Psi_{i}$ between $TN_{i}\oplus{\mathbb{R}}^{m+N-1}$ and $h_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$, satisfying the following conditions:
  • $N_{i}$ is a cobordism between $M_{i}$ and $M_{i+1}$.
  • The restrictions of $h_{i}$ to $M_{i}$ and $M_{i+1}$ coincide with $g_{i}$ and $g_{i+1}$, respectively.
  • The restrictions of $\Psi_{i}$ to $M_{i}$ and $M_{i+1}$ coincide with $\Phi_{i}$ and $\Phi_{i+1}$, respectively.
Here, regarding the identifications $TM_{i}\oplus{\mathbb{R}}\cong TN_{i}|_{M_{i}}$ and $TM_{i+1}\oplus{\mathbb{R}}\cong TN_{i}|_{M_{i+1}}$, the factor $\mathbb{R}$} corresponds to the inward-pointing normal vector on one boundary and the outward-pointing normal vector on the other.
We may assume that every connected component of $N_{i}$ intersects at least one of $M_{i}$ or $M_{i+1}$. This is because if a connected component of $N_{i}$ does not intersect either $M_{i}$ or $M_{i+1}$, that component can simply be removed from $N_{i}$.
First, we consider the case where $m\geq 2$. Since $N_{i}$ has a non-empty boundary by assumption, it is known that $N_{i}$ is homotopy equivalent to a CW complex of dimension $n$ or less by using a Morse function. Therefore, given the isomorphism $TN_{i}\oplus{\mathbb{R}}^{m+N-1}\stackrel{{\scriptstyle\Psi_{i}}}{{\cong}}h_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$, since $n+m\geq n+2$, Lemma 3.2 implies that there exists an isomorphism $\psi_{i}$ between $TN_{i}$ and $h_{i}^{*}\xi$, unique up to homotopy, such that $\Psi_{i}$ is homotopic to $\psi_{i}\oplus\mathrm{id}_{\mathbb{R}^{N}}$. Moreover, the isomorphisms between $TM_{i}\oplus\mathbb{R}^{m}$ and $g_{i}^{*}\xi$ induced by $\psi_{i-1}$ and $\psi_{i}$ on the boundary $M_{i}$ become homotopic to $\Phi_{i}$ after taking the direct sum with $\mathbb{R}^{N}$. Thus, the injectivity part of Lemma 3.2 implies that they are homotopic to each other. Consequently, these isomorphisms allow us to glue the cobordisms along the boundaries, so that the sequence $N_{0},N_{1},\ldots,N_{l}$ defines a zigzag in $\Omega^{\prime}(X;\xi,m)$, which proves that $M,g,\phi]=0\in\Omega^{\prime}(X;\xi,m)$.
Next, we show the case where $m=1$. The main difficulty in this case is determining whether the isomorphisms between $TN_{i}\oplus{\mathbb{R}}^{m-1}$ (which is $TN_{i}$) and $h_{i}^{*}\xi$ glue together along the boundaries. To resolve this issue, we proceed by considering the sequence starting from $N_{0}$. We first note that, as demonstrated in the $m\geq 2$ discussion, the isomorphism between $TN_{i}\oplus{\mathbb{R}}$ and $h_{i}^{*}\xi\oplus{\mathbb{R}}$ is uniquely determined up to homotopy.
Since the isomorphism between $TM_{0}\oplus{\mathbb{R}}$ and $g_{0}^{*}\xi$ is already fixed as $\phi$ over $M=M_{0}$, we need to construct an isomorphism between $TN_{0}$ and $h_{0}^{*}\xi$ over $N_{0}$ whose restriction to $M_{0}$ is compatible with $\phi$ via the isomorphism $TN_{0}\oplus{\mathbb{R}}\cong h_{0}^{*}\xi\oplus{\mathbb{R}}$.
This can be achieved by the following procedure:
  • First, we take a CW decomposition of $N_{0}$ as an $(n+1)$-dimensional CW complex.
  • Next, we construct an isomorphism between $TN_{0}$ and $h_{0}^{*}\xi$ over $N_{0}\setminus M_{0}$, proceeding sequentially from lower-dimensional cells. This construction involves extending a lift of the map $GL(n+1)\to GL(n+2)$ given on the boundary $\partial D^{r}$ to the entire cell $D^{r}$. As mentioned above, the obstruction to this lifting vanishes for dimensions $r\leq n$. Therefore, we can complete this construction up to the $n$-cells.
  • This leaves the $(n+1)$-cells. Using the isomorphism defined on $\partial D^{n+1}$, we can extend the isomorphism between $TN_{0}$ and $h_{0}^{*}\xi$ to $D^{n+1}$ minus a single interior point. This allows us to construct the desired isomorphism over $N_{0}$ excluding a finite number of points.
  • Finally, we excise the open neighborhoods of these finite points. This procedure yields a cobordism in $\Omega^{\prime}(X;\xi,1)$ between $M_{0}$ and $M_{1}\cup\left(\bigcup_{j=1}^{k_{1}}S^{n}\right)$.
By iteratively repeating this process, we find that $M$ and a finite union of $S^{n}$’s determine the same element in $\Omega^{\prime}(X;\xi,1)$. The remaining task is to determine whether these $S^{n}$ vanish in $\Omega^{\prime}(X;\xi,1)$.
This part constitutes the most technical aspect of the proof. The operation performed above implies that the remaining $S^{n}$ elements we must consider are of the form $S^{n},c:S^{n}\to X,\phi]\in\Omega^{\prime}(X;\xi,1)$, where:
  • The map $c:S^{n}\to X$ is a constant map.
  • When $\phi$ is viewed as a map $\phi:S^{n}\to GL(n+1)$ via the canonical isomorphism $TS^{n}\oplus{\mathbb{R}}\cong{\mathbb{R}}^{n+1}$, the class $\phi]$ maps to zero under the homomorphism $S^{n},GL(n+1)]\to[S^{n},GL(n+2)]$.
We now show that these elements vanish in $\Omega^{\prime}(X;\xi,m)$ by explicitly constructing a cobordism.
[…]
Next, consider the kernel of the map $\pi_{n}(GL^{+}(n+1))\to\pi_{n}(GL^{+}(n+2))$. Considering the long exact sequence of homotopy groups $\pi_{n+1}(S^{n+1})\cong{\mathbb{Z}}\to\pi_{n}(GL^{+}(n+1))\to\pi_{n}(GL^{+}(n+2))$, this kernel is a cyclic group generated by the clutching function $r\cdot c:S^{n}\to GL^{+}(n+1)$ associated with the trivializations of $TS^{n+1}$ on the northern and southern hemispheres, respectively.
Since $GL^{+}(n+1)$ is a path-connected H-space, all actions of the fundamental group on the higher homotopy groups are trivial. Thus, $S^{n},GL^{+}(n+1)]$ and $\pi_{n}(GL^{+}(n+1))$ can be naturally identified. Since $\phi$ maps to 0 under $S^{n},GL(n+1)]\to[S^{n},GL(n+2)]$, it follows that $\phi^{\prime}$ also maps to 0. Therefore, we can choose an integer $a$ such that $\phi^{\prime}]=a[r\cdot c]\in\pi_{n}(GL^{+}(n+1))$.
If $a=0$, this means that $\phi$ is null-homotopic, so $\phi$ can be directly extended to $D^{n+1}$. Thus, this case is null-cobordant.
Next, consider the case where $a$ is positive. In this case, we first decompose $S^{n}$ into $a$ copies of $S^{n}$ (more precisely, we construct a cobordism between $S^{n}$ and a disjoint union of $a$ copies of $S^{n}$) such that on each copy, the isomorphism $\phi$ is homotopic to either $r$ or $r\cdot c$. Here, geometrically considering the operation of multiplying the isomorphism $\phi$ by $r$ from the left, this corresponds to swapping the outward and inward directions of the corresponding $\mathbb{R}$} factor in the identification $TS^{n}\oplus{\mathbb{R}}\cong TD^{n+1}|_{S^{n}}$. Therefore, by swapping the outward/inward direction of $\oplus{\mathbb{R}}$, both $r$ and $r\cdot c$ are replaced by constant maps, which allows us to extend this $S^{n}$ to $D^{n+1}$. Thus, this case is also null-cobordant.
Here, when $n$ is even, the image of the homomorphism $\pi_{n+1}(GL(n+2))\to\pi_{n+1}(S^{n+1})\cong{\mathbb{Z}}$ induced by the continuous map $GL(n+2)\to S^{n+1}$ (taking the $(n+2)$-th column) contains $\chi(S^{n+2})=2$ [11]. Consequently, $2[r\cdot c]=0\in\pi_{n}(GL^{+}(n+1))$ holds. Therefore, $a$ can be taken as either 0 or $1$, which completes the proof for this case.
The remaining case is when $n$ is odd and $a$ is negative. First, similar to the positive case, we can reduce this to the case a=1a=-1. We begin by proving the following lemma.
Lemma 3.3.
For a continuous map $f:S^{n}\to GL(n+1)$, let $\chi_{f}$ denote the degree of the map $S^{n}\to S^{n}$ obtained by composing $f$ with the natural map $GL(n+1)\to S^{n}$. If $n$ is odd, then
$\chi_{r\cdot f}=2-\chi_{f}$
holds. In particular, $\chi_{r}=\chi_{r\cdot c}=2$.
Proof.
We construct a trivialization of $TS^{n+1}$ on the northern hemisphere of $S^{n+1}$ (excluding the north pole) from $f$, and a trivialization on the southern hemisphere (excluding the south pole) from $r\cdot f$. By the definition of $r$, these trivializations glue along the equator, providing a trivialization of $TS^{n+1}$ on $S^{n+1}$ excluding the north and south poles.
By extracting the $(n+1)$-th tangent vector of this trivialization, we obtain a vector field on $S^{n+1}$ (excluding the poles) that is nowhere zero. We apply the Poincaré-Hopf theorem to this vector field.
The indices at the north and south poles are $\chi_{f}$ and $\chi_{r\cdot f}$ by definition. Since the Euler characteristic of $S^{n+1}$ (which is even-dimensional for odd $n$) is 2, we obtain $\chi_{f}+\chi_{r\cdot f}=2$. This proves the lemma. ∎
Now, since $\phi^{\prime}]=-[r\cdot c]\in\pi_{n}(GL^{+}(n+1))$, we have $\chi_{\phi^{\prime}}=-\chi_{r\cdot c}=-2$. This implies $\chi_{r\cdot\phi^{\prime}}=2-(-2)=4$.
Therefore, by swapping the outward/inward direction of $\oplus{\mathbb{R}}$, the case a=1a=-1 reduces to the case $a=2$ (since the degree corresponds to $2[r\cdot c]$ which has degree 4). From the above, it is shown that all considered $S^{n}$ vanish in $\Omega^{\prime}(X;\xi,1)$. This shows that $\lambda$ is an isomorphism.
It remains to show the latter half of Theorem A; that is, the direct cobordism relation in $\Omega^{\prime}(X;\xi,m)$ is symmetric when $n\neq 1$. The strategy of this proof is analogous to the first half of the proof of Theorem 4.4 in [3].
Suppose $M_{0}\sim M_{1}$ via a cobordism $N$ between them. We consider constructing a vector field on $N$ that points outwards along the boundaries $M_{0}$ and $M_{1}$. If such a vector field is nowhere zero, we can compose the trivialization of $TN$ with the reflection defined by this vector field. This allows us to swap the inward and outward directions of the $\oplus{\mathbb{R}}$ factor at both $M_{0}$ and $M_{1}$, thereby establishing symmetry. If the Euler characteristic of $N$ is non-zero, such a vector field may not exist globally; however, one can always construct such a vector field away from a finite set of points. By excising open neighborhoods of these points from $N$, we obtain the relation $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k}S^{n}$.
It remains to show that these $S^{n}$ components are null-cobordant without switching the inward/outward direction of $\oplus{\mathbb{R}}$.
When $n$ is even, as previously stated, $2[r\cdot c]=0$. Thus, the element $\phi\in[S^{n},GL(n+1)]$ corresponding to the isomorphism is either null-homotopic or of the form $r\cdot(\text{null-homotopic element})$. In the null-homotopic case, the sphere clearly bounds $D^{n+1}$. In the case of $r\cdot(\text{null-homotopic element})$, we previously noted that the sphere bounds $D^{n+1}$ if we switch the boundary direction. This implies that if we take a parallelizable closed $(n+1)$-dimensional manifold and remove $D^{n+1}$, the remaining manifold bounds the $S^{n}$ in question. For example, considering the torus $T^{n+1}=S^{1}\times\cdots\times S^{1}$, we see that $T^{n+1}\setminus\mathrm{int}(D^{n+1})$ serves as a bounding manifold.
Next, consider the case where $n$ is an odd integer with $n\geq 3$. As discussed before, for the element $\phi$ corresponding to the isomorphism, we may assume $\phi^{\prime}]$ is one of $0,[r\cdot c]$, or $r\cdot c]$ in $\pi_{n}(GL^{+}(n+1))$. The case 0 is bounded by $D^{n+1}$, and the case $r\cdot c]$ is bounded by $T^{n+1}\setminus\mathrm{int}(D^{n+1})$ as described above. It remains to find an $(n+1)$-dimensional compact manifold that bounds the case $r\cdot c]$. To state the conclusion, we can use $S^{\frac{n+1}{2}}\times S^{\frac{n+1}{2}}\setminus\mathrm{int}(D^{n+1})$ when $n+1$ is a multiple of 4, and $S^{\frac{n-1}{2}}\times S^{\frac{n+3}{2}}\setminus\mathrm{int}(D^{n+1})$ when n1n-1 is a multiple of 4.
We demonstrate this below. Let $X^{n+1}$ denote $S^{\frac{n+1}{2}}\times S^{\frac{n+1}{2}}$ if $n+1$ is divisible by 4, and $S^{\frac{n-1}{2}}\times S^{\frac{n+3}{2}}$ if n1n-1 is divisible by 4. Note that $X^{n+1}$ is a stably trivial, connected, closed $(n+1)$-dimensional manifold with Euler characteristic $4$ (this proof requires $n\neq 1$, as connectedness fails otherwise). Since a stably trivial closed manifold is almost parallelizable, $TX^{n+1}$ admits a trivialization on $X^{n}$ minus a single point.
Consider the obstruction to extending this trivialization to the point. This obstruction lies in $S^{n},GL(n+1)]$. Since $X^{n+1}$ is stably trivial, this element maps to 0 under the stabilization map $S^{n},GL(n+1)]\to[S^{n},GL(n+2)]$.
Applying the Poincaré-Hopf theorem to $X^{n+1}$, the image of this element under the map $S^{n},GL(n+1)]\to[S^{n},S^{n}]$ coincides with the Euler characteristic of $X^{n+1}$, which is $4$. By the lemma stated earlier, swapping the outward and inward directions of $\oplus{\mathbb{R}}$ transforms an element of degree $4$ into an element of degree 24=22-4=-2 in $S^{n},GL(n+1)]$. Consequently, an $S^{n}$ carrying an element of degree 2-2 (corresponding to $r\cdot c]$ since the degree of $r\cdot c]$ is $2$) is bounded by $X^{n+1}\setminus\mathrm{int}(D^{n+1})$.
In conclusion, all such $S^{n}$ are null-cobordant without requiring a change in the boundary normal direction. Thus, $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k_{1}}S^{n}\sim M_{0}$, proving the symmetric property.
This completes the proof of Theorem A.
Remark 3.4.
When $n=1$, the symmetric property does not necessarily hold. See Example 4.6 in [3].
Remark 3.5.
We present a simple consequence of this theorem. By [7, Theorem 3.0.8], the group $\pi_{1}(MTSO(2))\cong\pi_{-1}(BSO(2)^{-(L_{2}^{SO}\oplus{\mathbb{R}}^{2})})\cong\Omega^{\prime}(BSO(2);L^{SO}_{2}\oplus{\mathbb{R}}^{2},1)$ vanishes. Since the cobordism group $\Omega^{\prime}(BSO(2);L^{SO}_{2}\oplus{\mathbb{R}}^{2},1)$ is trivial, every element is null-cobordant (i.e., it bounds some manifold).
In particular, given a pair $(M,X)$ consisting of a closed oriented $3$-manifold $M$ and a nowhere-vanishing vector field $X$ on $M$, there exists a triple $(N,Y_{1},Y_{2})$ consisting of a compact oriented $4$-manifold NN and two linearly independent vector fields $Y_{1},Y_{2}$ on $N$ such that:
  • $\partial N\cong M$;
  • $Y_{1}|_{M}=X$, and $Y_{2}|_{M}$ is the outward-pointing normal vector field.
This result is stronger than the fact that the $3$-dimensional oriented cobordism group is trivial; it means that we can impose additional structure on the bounding manifolds.
4. The proof of Theorem B
In this section, we provide the proofs of Theorem B1, B2, and B3. All proofs rely on the methods detailed in Section 5 of [5]. The key result we employ is the following theorem, which describes a fibrational relationship between Thom spectra.
Theorem 4.1 (Theorem 5.1, [5]).
Let $X$ be a topological space, $V$ a virtual vector bundle over $X$, and $W$ a real vector bundle over $X$. Let $f:S(W)\to X$ be the spherical fibration of $W$. Then, the following is a fibration of spectra:
$S(W)^{f^{\ast}V}\to X^{V}\to X^{V\oplus W}$
Proof of Theorem B.
First, we show that the vertical columns of the $3\times 3$ diagrams are fibrations of spectra. This follows from the fact that the projection of the sphere bundle $S(L_{n+1}^{O})$ of the universal rank $(n+1)$ vector bundle $L_{n+1}^{O}\to BO(n+1)$ can be identified with the map $BO(n)\to BO(n+1)$ (via the inclusion $O(n)\hookrightarrow O(n+1)$) due to the identification:
$S(L_{n+1}^{O})=EO(n+1)\times_{O(n+1)}S^{n}=EO(n+1)\times_{O(n+1)}O(n+1)/O(n)\cong EO(n+1)/O(n).$
The same reasoning applies to the spherical bundles corresponding to the maps $B\mathrm{Spin}(n)\to B\mathrm{Spin}(n+1)$, $B\mathrm{Pin}^{-}(n)\to B\mathrm{Pin}^{-}(n+1)$, and $B\mathrm{Spin}^{c}(n)\to B\mathrm{Spin}^{c}(n+1)$. Thus, applying Theorem 5.1 from [5] to these spherical fibrations shows that all vertical columns in the $3\times 3$ diagrams of Theorem B are fibrations of spectra.
Next, we explain why the horizontal rows of the $3\times 3$ diagrams are fibrations of spectra. First, for the case of $BO(n)$, we consider the spherical bundle of the determinant line bundle $\mathrm{det}L_{n}^{O}$. We have the identification:
$S(\mathrm{det}L_{n}^{O})=EO(n)\times_{O(n)}\{\pm 1\}=EO(n)\times_{O(n)}O(n)/SO(n)\cong EO(n)/SO(n).$
This spherical bundle is the natural map $BSO(n)\to BO(n)$ induced by the inclusion $SO(n)\hookrightarrow O(n)$. Thus, applying Theorem 5.1 to $W=\mathrm{det}L_{n}^{O}$ shows that the horizontal rows of Theorem B1 are fibrations of spectra.
Similarly, since $\mathrm{Pin}^{-}(n)/\mathrm{Spin}(n)\cong\{\pm 1\}$, the same argument shows that the horizontal rows of Theorem B2 are also fibrations of spectra.
Finally, the natural inclusion $\mathrm{Spin}(n)\to\mathrm{Spin}^{c}(n)$ induces a Lie group fibration $\mathrm{Spin}(n)\to\mathrm{Spin}^{c}(n)\to S^{1}$. Since $\mathcal{O}_{n}$ is the pullback of the canonical complex line bundle over $BS^{1}$ via the map $B\mathrm{Spin}^{c}(n)\to BS^{1}$, and we have $\mathrm{Spin}^{c}(n)/\mathrm{Spin}(n)\cong S^{1}$, the spherical bundle $S(\mathcal{O}_{n})$ is the map $B\mathrm{Spin}(n)\to B\mathrm{Spin}^{c}(n)$. Therefore, the horizontal rows of Theorem B3 also give fibrations of spectra. This completes the proofs of Theorems B1, B2, and B3.
5. The proof of theorem C
In this section, we provide the proofs of Theorems C1, C2, and C3. For the first part of each proof, we utilize Theorem A. For the second part, we employ the long exact sequences of homotopy groups induced by the $3\times 3$ diagrams established in Theorem B.
Proof of Theorem C1.
Based on Theorem A, we examine the geometric structure of the cobordism group corresponding to $\Omega^{\prime}(BO(n);L^{O}_{n}\oplus\mathrm{det}L^{O}_{n},1)\cong\pi_{-1}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})})$.
According to Theorem A, each element of this cobordism group is represented by a triple: an $n$-dimensional closed manifold $M$, a map $g:M\to BO(n)$, and a stable isomorphism $\phi:TM\oplus{\mathbb{R}}\cong g^{*}(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})$. First, observe that the bundle $L^{O}_{n}\oplus\mathrm{det}L^{O}_{n}$ carries a natural orientation. Consequently, this induces a natural orientation on $M$ via the isomorphism $\phi$. Furthermore, via the isomorphism $\phi$, we can uniquely identify a line subbundle $\eta$ of $TM\oplus{\mathbb{R}}$ that corresponds to the factor $g^{\ast}\mathrm{det}L^{O}_{n}$.
Conversely, suppose we are given a pair $(M,\eta)$ consisting of a closed oriented $n$-manifold $M$ and a line subbundle $\eta$ of $TM\oplus{\mathbb{R}}$. Let $\eta^{\perp}$ denote the orthogonal complement of $\eta$ in $TM\oplus{\mathbb{R}}$, and let $g:M\to BO(n)$ be a classifying map for $\eta^{\perp}$. Since the direct sum $\eta^{\perp}\oplus\eta=TM\oplus{\mathbb{R}}$ is an oriented vector bundle (as $M$ is oriented), the line bundle $\eta$ acts as the determinant bundle of $\eta^{\perp}$. Thus, there exists a natural isomorphism between $\eta$ and $\mathrm{det}(\eta^{\perp})$. Using this identification, we obtain a natural isomorphism between $TM\oplus{\mathbb{R}}=\eta^{\perp}\oplus\eta$ and the pullback bundle $g^{*}(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})$.
Since these correspondences are inverse to each other, the first part of Theorem C1 is established.
It remains to compute the group $\pi_{-1}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})})$. We explain this computation using the diagram provided in Theorem B1.
First, to compute the homotopy groups appearing in the diagram, we utilize the following known lemmas.
Lemma 5.1.
Let $V\to X$ be a virtual vector bundle of positive rank. Then $\pi_{0}(X^{V})=0$.
Lemma 5.2.
Let $X$ be a path-connected space, and let $V\to X$ be a virtual vector bundle of rank 0. Then
$\pi_{0}(X^{V})\cong\begin{cases}{\mathbb{Z}},&\text{if }V\text{ is orientable}\\ {\mathbb{Z}}/2{\mathbb{Z}},&\text{if }V\text{ is non-orientable}\end{cases}$
These two lemmas follow immediately from the Adams spectral sequence.
Lemma 5.3.
Let $n$ be a non-negative integer and $m$ be a positive integer. Then $\pi_{-m}(BO(n)^{-L_{n}^{O}})$ is isomorphic to the (nm)(n-m)-dimensional cobordism group, $\Omega_{n-m}^{O}$. A similar theorem holds for $BSO$, $B\mathrm{Spin}$, and $B\mathrm{Pin}^{-}$. Here, for convenience, we define $\Omega_{n-m}^{O}=0$ when $n<m$.
Proof.
Consider the fibration of spectra:
$BO(n)^{-(L_{n}^{O}\oplus{\mathbb{R}})}\to BO(n+1)^{-L_{n+1}^{O}}\to BO(n+1)^{0}$
This induces the following long exact sequence of homotopy groups:
$\pi_{-m}(BO(n+1)^{0})\to\pi_{-m}(BO(n)^{-L_{n}^{O}})\to\pi_{-(m+1)}(BO(n+1)^{-L_{n+1}^{O}})\to\pi_{-(m+1)}(BO(n+1)^{0})$
The first and last terms are zero by Lemma 5.1. Therefore, we obtain the isomorphism
$\pi_{-m}(BO(n)^{-L_{n}^{O}})\cong\pi_{-(m+1)}(BO(n+1)^{-L_{n+1}^{O}}).$
By iterating this process, we get
$\pi_{-m}(BO(n)^{-L_{n}^{O}})=\pi_{n-m}(BO(n)^{{\mathbb{R}}^{n}-L_{n}^{O}})\cong\pi_{n-m}(MTO)\cong\Omega^{O}_{n-m}.$
This completes the proof. ∎
Remark 5.4.
It is known, due to Ebert [6], that $\pi_{0}(BO(n)^{-L_{n}^{O}})$ and $\pi_{0}(BSO(n)^{-L_{n}^{SO}})$ are isomorphic to the cobordism groups defined by Reinhart in [9]. Furthermore, [2] provides computations of $\pi_{m}(BSO(n)^{-L_{n}^{SO}})$ for $0\leq m\leq 3$ using the Adams spectral sequence.
Using these results, by examining the homotopy groups of degrees 2-2 and 1-1 in Theorem B1, we construct the following commutative diagram in which every row and column is exact.
[…]
First, we consider the second row from the bottom. The map $\Omega_{n}^{SO}\to\pi_{-2}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})$ is induced by the bundle map $L^{SO}_{n+1}\oplus{\mathbb{R}}\to L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1}$. Therefore, an element $M]\in\Omega_{n}^{SO}$ maps to the class $M,g:M\to BO(n+1),TM\oplus{\mathbb{R}}^{2}\stackrel{{\scriptstyle\phi}}{{\cong}}g^{\ast}(L_{n+1}^{O}\oplus\mathrm{det}L_{n+1}^{O})]$. Here, $g$ is the classifying map for $TM\oplus{\mathbb{R}}$, and $\phi$ is the combination of the two natural isomorphisms $TM\oplus{\mathbb{R}}\cong g^{\ast}L_{n+1}^{O}$ and ${\mathbb{R}}\cong g^{\ast}\mathrm{det}L_{n+1}^{O}$. This map admits a natural left inverse $q:[M,g,\phi]\mapsto[M]$.
By the same argument, the map $\Omega_{n-1}^{SO}\to\pi_{-3}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})$ is also injective. Thus, by considering the exact sequence
$\pi_{-2}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})})$$\Omega_{n-1}^{O}$$\Omega_{n-1}^{SO}$$\pi_{-3}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})$$\scriptstyle{r}$$\scriptstyle{0}$
we deduce the surjectivity of $r$.
Therefore, we find that the second row from the bottom splits, yielding the isomorphism
$r\oplus q:\pi_{-2}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})\xrightarrow{\cong}\Omega^{O}_{n-1}\oplus\Omega^{SO}_{n}.$
By composing this isomorphism with $\pi_{-1}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})})\to\pi_{-2}(BO(n)^{-(L^{O}_{n}\oplus\mathrm{det}L^{O}_{n})})$ , we obtain the map $\epsilon$.
From the construction of $q$, the second component of this map is clearly $M,\eta]\mapsto[M]$. Regarding the first component $r^{\prime}$, it is known from [5, Proposition 3.17] that for a class $M,\eta]$, this map corresponds to taking a generic section $s$ of $\eta$ transverse to the zero section and returning the class $s^{-1}(0)]$.
Thus, it remains to determine the kernel of this surjective group homomorphism $\epsilon$. To achieve this, it suffices to determine whether the map $p:\pi_{-1}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map or surjective.
Again, by [5, Proposition 3.17], $p$ is the map that assigns to a class $M,\eta]$ the parity of the cardinality of $t^{-1}(0)$, where $t$ is a generic section of $\eta^{\perp}$ transverse to the zero section. Since the dimension of $M$ and the rank of $\eta^{\perp}$ are both $n+1$, this value coincides with the pairing $\langle w_{n+1}(\eta^{\perp}),[M]\rangle$.
Here, by the properties of Stiefel-Whitney classes, the equality
$w_{n+1}(\eta^{\perp})=w_{n}(\eta^{\perp})w_{1}(\eta)+w_{n+1}(TM\oplus{\mathbb{R}})$
holds. On the right-hand side, the term $w_{n+1}(TM)$ corresponds to the parity of the Euler characteristic of $M$. We now consider the term $w_{n}(\eta^{\perp})w_{1}(\eta)$. Let $s$ be a generic section of $\eta$ transverse to the zero section, and consider the zero locus $Z=s^{-1}(0)\subset M$. By construction, there is a canonical isomorphism between the normal bundle of $Z$ and the restriction $\eta|_{Z}$. Consequently, we have $TZ\oplus{\mathbb{R}}\cong\eta^{\perp}|_{Z}$. Under the Gysin (Umkehr) homomorphism $H^{n}(Z;{\mathbb{Z}}/2{\mathbb{Z}})\to H^{n+1}(M;{\mathbb{Z}}/2{\mathbb{Z}})$, the class $w_{n}(\eta^{\perp}|_{Z})=w_{n}(TZ)$ maps to $w_{n}(\eta^{\perp})w_{1}(\eta)$. Therefore, $w_{n}(\eta^{\perp})w_{1}(\eta)$ corresponds to the parity of the Euler characteristic of $Z$ (which is $s^{-1}(0)$).
From the above, it follows that $p$ coincides with the composition of the map $\epsilon:\pi_{-1}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})\to\Omega^{O}_{n}\oplus\Omega^{SO}_{n+1}$ and the map that takes the sum of the parities of the Euler characteristics of the two components.
Regarding closed manifolds:
  • For unoriented manifolds (corresponding to $\Omega^{O}_{n}$), the Euler characteristic is always 0 in odd dimensions due to the orientable double cover and Poincaré duality. In even dimensions, it can be odd, as seen by considering $\mathbb{RP}^{2n}$.
  • For oriented closed manifolds (corresponding to $\Omega^{SO}_{n+1}$), the Euler characteristic is always even in odd dimensions due to Poincaré duality. Furthermore, in dimensions congruent to $2\pmod{4}$, the intersection form in the middle dimension is skew-symmetric, which implies the middle Betti number is even, so the Euler characteristic is also even. On the other hand, considering $\mathbb{CP}^{2n}$, we see that the Euler characteristic can be odd in dimensions divisible by 4.
Combining these facts, We see the map $p:\pi_{-1}(BO(n+1)^{-(L^{O}_{n+1}\oplus\mathrm{det}L^{O}_{n+1})})\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map only when $n\equiv 1\pmod{4}$, and is surjective in all other cases.
This concludes the proof of Theorem C1. ∎
Proof of Theorem C2.
The proof proceeds in a manner analogous to that of Theorem C1.
First, as indicated by the following fibration:
[…]
the condition for a vector bundle $E$ to admit a $\mathrm{Pin}^{-}$ structure is $w_{2}(E)+w_{1}(E)^{2}=0$. Therefore, since
$w_{2}(L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n})=w_{2}(L^{\mathrm{Pin}^{-}}_{n})+w_{1}(L^{\mathrm{Pin}^{-}}_{n})w_{1}(\mathrm{det}L^{\mathrm{Pin}^{-}}_{n})=w_{2}(L^{\mathrm{Pin}^{-}}_{n})+w_{1}(L^{\mathrm{Pin}^{-}}_{n})^{2}=0,$
the bundle $L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n}$ admits a $\mathrm{Spin}$ structure. We fix this $\mathrm{Spin}$ structure.
By Theorem A, each element of $\Omega^{\prime}(B\mathrm{Pin}^{-}(n);L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n},1)$ is represented by three components: an $n$-dimensional closed manifold $M$, a map $g:M\to B\mathrm{Pin}^{-}(n)$, and a stable isomorphism $\phi:TM\oplus{\mathbb{R}}\cong g^{*}(L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n})$. The fixed $\mathrm{Spin}$ structure on the target bundle induces a $\mathrm{Spin}$ structure on $M$ via the isomorphism $\phi$. Furthermore, via the isomorphism $\phi$, we can identify a line subbundle $\eta$ of $TM\oplus{\mathbb{R}}$ corresponding to $\mathrm{det}L^{\mathrm{Pin}^{-}}_{n}$.
Conversely, given a pair $(M,\eta)$ of an $n$-dimensional closed $\mathrm{Spin}$ manifold $M$ and a line subbundle $\eta$ of $TM\oplus{\mathbb{R}}$, we have $w_{2}(\eta^{\perp}\oplus\eta)=0$, which implies $w_{2}(\eta^{\perp})=w_{1}(\eta^{\perp})w_{1}(\eta)=w_{1}(\eta^{\perp})^{2}$. Thus, $\eta^{\perp}$ admits a $\mathrm{Pin}^{-}$ structure. Although there may be multiple choices for this $\mathrm{Pin}^{-}$ structure, we define $g:M\to B\mathrm{Pin}^{-}(n)$ and $\phi$ such that the bundle map from $TM\oplus{\mathbb{R}}=\eta^{\perp}\oplus\eta$ to $L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n}$ preserves the $\mathrm{Spin}$ structure.
Since these correspondences are inverses of each other, the first part of Theorem C2 is established.
The second part of the proof follows the same logic as in Theorem C1, with $SO$ replaced by $\mathrm{Spin}$ and $O$ replaced by $\mathrm{Pin}^{-}$, up to the final step. It remains to determine the values of $n$ for which the map $p:\pi_{-1}(B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1})})\to\Omega^{\mathrm{Pin}^{-}}_{n}\oplus\Omega^{\mathrm{Spin}}_{n+1}\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map.
  • For closed $\mathrm{Pin}^{-}$ manifolds, the Euler characteristic is always 0 in odd dimensions by the orientable double cover and Poincaré duality. Furthermore, according to [8, Table 2.2], it is known that the Euler characteristic is always even in dimensions congruent to $4\pmod{8}$. On the other hand, considering $\mathbb{HP}^{2n}$ and $\mathbb{RP}^{4n+2}$, the Euler characteristic can be odd in dimensions congruent to $0,2,6\pmod{8}$.
  • For closed $\mathrm{Spin}$ manifolds, as mentioned earlier, the Euler characteristic is always even in odd dimensions and in dimensions congruent to $2\pmod{4}$. When the dimension is congruent to $4\pmod{8}$, the parity of the Euler characteristic coincides with that of the middle Betti number, which in turn coincides with the parity of the signature (due to the intersection form). By Ochanine’s theorem, the signature of a Spin manifold of dimension $4\pmod{8}$ is always divisible by 16; thus, the Euler characteristic must be even. On the other hand, considering $\mathbb{HP}^{2n}$, the Euler characteristic can be odd in dimensions divisible by 8.
Combining these facts, the map $p$ is the zero map only when $n\equiv 1,3,4,5\pmod{8}$, and is surjective otherwise. This concludes the proof of Theorem C2. ∎
Proof of Theorem C3.
The proof proceeds in a manner analogous to that of Theorem C1.
First, note that $L^{\mathrm{Spin}^{c}}_{n}$ and $\mathcal{O}_{n}$ are both oriented vector bundles over $B\mathrm{Spin}^{c}(n)$. They are defined as the pullbacks of the canonical oriented vector bundle and the canonical complex line bundle via the maps
$B\mathrm{Spin}^{c}(n)\to BSO(n)\times BU(1)\xrightarrow{pr_{1}}BSO(n)$
and
$B\mathrm{Spin}^{c}(n)\to BSO(n)\times BU(1)\xrightarrow{pr_{2}}BU(1),$
respectively. Since the map $B\mathrm{Spin}^{c}(n)\to BSO(n)\times BU(1)$ pulls back the non-zero elements of $H^{2}(BSO(n);{\mathbb{Z}}/2{\mathbb{Z}})\cong{\mathbb{Z}}/2{\mathbb{Z}}$ and $H^{2}(BU(1);{\mathbb{Z}}/2{\mathbb{Z}})\cong{\mathbb{Z}}/2{\mathbb{Z}}$ to the same element, we have
$w_{2}(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})=w_{2}(L^{\mathrm{Spin}^{c}}_{n})+w_{2}(\mathcal{O}_{n})=0.$
Thus, the direct sum $L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n}$ admits a $\mathrm{Spin}$ structure. We fix this $\mathrm{Spin}$ structure.
By Theorem A, each element of $\Omega^{\prime}(B\textnormal{Spin}^{c}(n);L^{\textnormal{Spin}^{c}}_{n}\oplus\mathcal{O}_{n},2)$ is represented by three components: an $n$-dimensional closed manifold $M$, a map $g:M\to B\mathrm{Spin}^{c}(n)$, and a stable isomorphism $\phi:TM\oplus{\mathbb{R}}^{2}\cong g^{*}(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})$. The fixed $\mathrm{Spin}$ structure on the target bundle induces a $\mathrm{Spin}$ structure on $M$ via the isomorphism $\phi$. Furthermore, via the isomorphism $\phi$, we can identify a rank 2 oriented subbundle $\eta$ of $TM\oplus{\mathbb{R}}^{2}$ corresponding to $\mathcal{O}_{n}$.
Conversely, given a pair $(M,\eta)$ of an $n$-dimensional closed $\mathrm{Spin}$ manifold $M$ and a rank 2 subbundle $\eta$ of $TM\oplus{\mathbb{R}}^{2}$, we have $w_{2}(\eta^{\perp}\oplus\eta)=0$, which implies $w_{2}(\eta^{\perp})=w_{2}(\eta)$. This condition allows $\eta$ to admit a $\mathrm{Spin}^{c}$ structure. Although there may be multiple choices for this $\mathrm{Spin}^{c}$ structure, we define $g:M\to B\mathrm{Spin}^{c}(n)$ and $\phi$ such that the bundle map from $TM\oplus{\mathbb{R}}^{2}=\eta^{\perp}\oplus\eta$ to $L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n}$ preserves the $\mathrm{Spin}$ structure.
Since these correspondences are inverses of each other, the first part of Theorem C3 is established.
Next, we show the second part. By taking the homotopy groups of degrees 3-3 and 2-2 in Theorem B3, we construct the following commutative diagram where every row and column is exact.
[…]
Based on the commutative diagram constructed above, we now proceed to extract information about the group $\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$.
As in the proof of Theorem C1, the map $\Omega_{n}^{\mathrm{Spin}}\to\pi_{-3}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})$ admits a natural left inverse $q:[M,g,\phi]\mapsto[M]$. The surjectivity of $r$ follows similarly.
Therefore, the second row from the bottom splits, yielding the isomorphism
$r\oplus q:\pi_{-3}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})\xrightarrow{\cong}\Omega^{\mathrm{Spin}^{c}}_{n-2}\oplus\Omega^{\mathrm{Spin}}_{n}.$
By composing this with the natural map $\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})\to\pi_{-3}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})$, we obtain the map $\epsilon$. From the construction of $q$, the second component of this map is clearly $M,\eta]\mapsto[M]$. Regarding the first component $r^{\prime}$, as mentioned before, it corresponds to taking a generic section $s$ of $\eta$ transverse to the zero section and returning $s^{-1}(0)]$.
It remains to determine the kernel of this surjective group homomorphism $\epsilon$. For this purpose, it suffices to determine the image of the map
$p:\pi_{-2}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})\to{\mathbb{Z}}.$
Again by [5, Proposition 3.17], $p$ is the map that assigns to a class $M,\eta]$ the number of points in $t^{-1}(0)$, where $t$ is a generic section of $\eta^{\perp}$ transverse to the zero section. Since the dimension of $M$ and the rank of $\eta^{\perp}$ are both $n+1$, and $\eta$ (and thus $\eta^{\perp}$) is oriented, this coincides with the pairing $\langle e(\eta^{\perp}),[M]\rangle$.
Since $2e=0$ in odd dimensions (for the vector bundle, i.e., when $n+1$ is odd), $p$ is the zero map when $n$ is even. When $n$ is odd, considering the element $S^{n+1},{\mathbb{R}}^{2}\subset TS^{n+1}\oplus{\mathbb{R}}^{2}]$, its image under $p$ is the Euler characteristic of $S^{n+1}$, which is $2$. Thus, $2$ is contained in the image of $p$. It remains only to consider the parity of the values of $p$.
The mod 2 reduction of the Euler class $e$ is $w_{n+1}$, and we have
$w_{n+1}(\eta^{\perp})=w_{n-1}(\eta^{\perp})w_{2}(\eta)+w_{n+1}(TM\oplus{\mathbb{R}}^{2}).$
On the right-hand side, $w_{n+1}(TM\oplus{\mathbb{R}}^{2})=w_{n+1}(TM)$ corresponds to the parity of the Euler characteristic of $M$. Regarding $w_{n-1}(\eta^{\perp})w_{2}(\eta)$, similar to Theorem C1, if we take a generic section $s$ of $\eta$ transverse to the zero section, this term corresponds to the parity of the Euler characteristic of $s^{-1}(0)$.
Thus, the mod 2 reduction of $p$ coincides with the composition of $\epsilon:\pi_{-2}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})\to\Omega^{\mathrm{Spin}^{c}}_{n-1}\oplus\Omega^{\mathrm{Spin}}_{n+1}$ and the map that takes the sum of the parities of the Euler characteristics. As mentioned earlier, the Euler characteristic of a closed Spin manifold can be odd only in dimensions divisible by 8. Regarding closed $\mathrm{Spin}^{c}$ manifolds, since they are oriented, the Euler characteristic can be odd only in dimensions divisible by 4. Indeed, $\mathbb{CP}^{2m}$ admits a $\mathrm{Spin}^{c}$ structure (as it is complex) and has an odd Euler characteristic.
Based on the above, the map $p:\pi_{-2}(B\mathrm{Spin}^{c}(n+1)^{-(L^{\mathrm{Spin}^{c}}_{n+1}\oplus\mathcal{O}_{n+1})})\to{\mathbb{Z}}$ behaves as follows:
  • If $n$ is even, it is the zero map.
[…]
  • If $n\equiv 1,5,7\pmod{8}$, it is surjective.
This concludes the proof of Theorem C3. ∎
6. Application
First, as a consequence of Theorem C2, we obtain the following result.
Theorem 6.1.
Consider the following condition for a non-negative integer $n$: ”There exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact $\mathrm{Spin}$ manifold $M$ with boundary $\partial M=S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta|_{S^{n}}=(TS^{n})^{\perp}$ (i.e., $\eta$ is orthogonal to the boundary along $S^{n}$).”
Then, this condition holds if and only if $n\equiv 0,2,6,7\pmod{8}$.
Proof.
This proposition holds if and only if the element $S^{n},{\mathbb{R}}\subset TS^{n}\oplus{\mathbb{R}}]$ vanishes in $\Omega^{\textnormal{Spin, line}}_{n}$. Combining [6, Prop. A.5] with the commutative diagram used in the proof of Theorem C2, we observe that the image of $1\in{\mathbb{Z}}/2{\mathbb{Z}}$ under the map $s:{\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{-1}(B\mathrm{Pin}^{-}(n)^{-(L^{\mathrm{Pin}^{-}}_{n}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n})})$ coincides with $S^{n},{\mathbb{R}}\subset TS^{n}\oplus{\mathbb{R}}]$. Therefore, the vanishing of this element is equivalent to the surjectivity of the map $p$ (discussed in the proof of Theorem C2). As established in the proof of Theorem C2, this occurs if and only if $n\equiv 0,2,6,7\pmod{8}$. This concludes the proof. ∎
Remark 6.2.
If we remove the $\mathrm{Spin}$ constraint, we can take $M=\mathbb{RP}^{n+1}\setminus\mathrm{int}(D^{n+1})$. Indeed, consider the compact manifold obtained by removing open neighborhoods of the north and south poles from $S^{n+1}$, equipped with a nowhere-vanishing vector field pointing from the north pole to the south pole. By identifying antipodal points, the vectors at identified points differ by a factor of 1-1. This allows us to construct a line subbundle of the tangent bundle on $\mathbb{RP}^{n+1}\setminus\mathrm{int}(D^{n+1})$ that is orthogonal to the boundary.
Characteristic class computations show that $\mathbb{RP}^{n+1}$ is oriented when $n$ is even, and admits a $\mathrm{Spin}$ structure when $n\equiv 2\pmod{4}$. Thus, for $n\equiv 2,6\pmod{8}$, this provides the desired construction.
When $n$ is odd, according to [4], the given condition is equivalent to the Euler characteristic of $M$ being zero. Consequently, the proposition is equivalent to the existence of a closed $(n+1)$-dimensional $\mathrm{Spin}$ manifold with Euler characteristic $1$. As mentioned previously, such manifolds exist only when $n\equiv 7\pmod{8}$ (i.e., when the dimension $n+1$ is divisible by 8), in which case $M=\mathbb{HP}^{\frac{n+1}{4}}\setminus\mathrm{int}(D^{n+1})$ satisfies the condition.
Clearly, for $n=0$, we can simply take $M=D^{1}$. On the other hand, the author is not aware of how to construct such an $M$ when $n$ is a multiple of 8 other than 0, nor is there an alternative proof known to the author explaining why such an $M$ does not exist when $n\equiv 4\pmod{8}$. Constraints on line fields are detailed in [4], but for even $n$, strong constraints have not yet been obtained, and the problem remains open.
Finally, we conclude this paper by proving the result of [1] using Theorem C3.
First, the cobordism group introduced in [1] can be redefined in the context of this paper as follows.
Definition 6.3.
Let $n$ be a positive integer. We consider the set of $n$-dimensional closed $\mathrm{Spin}$ manifolds $M$ such that the Euler characteristic of every connected component is 0. We write $M_{1}\sim M_{2}$ if there exists a pair $(W,\xi)$ satisfying the following conditions:
  • $W$ is a compact $(n+1)$-dimensional $\mathrm{Spin}$ manifold with boundary $\partial W=M_{1}\cup M_{2}$.
  • $\xi$ is a rank 2 oriented subbundle of $TW$.
  • Restricted to $M_{i}$, $\xi$ decomposes as the direct sum of the trivial line bundle $\mathbb{R}$} (which is orthogonal to the boundary) and a trivial line subbundle $\eta$ of $TM_{i}$.
We denote by $\Omega_{2,n-1}^{\mathrm{Spin}_{0}}$ the cobordism group defined by taking the quotient by this equivalence relation.
In the following, we rephrase this equivalence relation in a simpler form.
First, let $M_{1}$ and $M_{2}$ be any two closed $\mathrm{Spin}$ manifolds such that the Euler characteristic of every connected component is 0. We prove the following lemma.
Lemma 6.4.
The relation $M_{1}\sim M_{2}$ holds if and only if the images of $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ and $M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]$ in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ coincide under the map
$\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})\to\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$
obtained by taking the homotopy groups (with appropriate degree shifts) of the commutative diagram in Theorem B3.
Proof.
First, we show necessity. Assume $M_{1}\sim M_{2}$. By the definition of the equivalence relation, we can choose trivial line subbundles $\eta_{1}\subset TM_{1}$ and $\eta_{2}\subset TM_{2}$ such that the classes
$M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]\quad\text{and}\quad[M_{2},\eta_{2}\oplus{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]$
define the same element in $\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$.
Consider the following commutative diagram obtained from Theorem B3:
${\mathbb{Z}}/2{\mathbb{Z}}$${\mathbb{Z}}/2{\mathbb{Z}}$$\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$$\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$$\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$$\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$}
In the second commutative square from the top, consider the images of $M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ and $M_{2},\eta_{2}\oplus{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]$. Since they map to the same element in the horizontal direction (to the right), they must map to the same element in the composition
$\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})\to\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})\to\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})}).$
The map $\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})\to\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ corresponds to forgetting one dimension of a rank 2 trivial subbundle (stabilization). Thus, the classes $M_{i},{\mathbb{R}}\subset TM_{i}\oplus{\mathbb{R}}]$ map to the same element in the final group. This proves necessity.
Next, we show sufficiency. Since the Euler characteristic of every connected component of $M_{1}$ and $M_{2}$ is 0, we can choose trivial line subbundles $\eta_{i}$ of $TM_{i}$. By assumption, the images of $M_{i},\eta_{i}\oplus{\mathbb{R}}\subset TM_{i}\oplus{\mathbb{R}}]$ coincide when mapped to $\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$. Consequently, the difference
$M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]-[M_{2},\eta_{2}\oplus{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]$
in $\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$ lies in the image of the map from ${\mathbb{Z}}/2{\mathbb{Z}}$.
If this element is 0, then $M_{1}\sim M_{2}$ holds by the definition of the equivalence relation. If this element is the image of the non-zero element of ${\mathbb{Z}}/2{\mathbb{Z}}$, we can adjust it using the image of $1$ under the map ${\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$. Specifically, we can replace $M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ with another element $M_{1},\eta^{\prime}_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ such that it maps to the same element $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ (the base class remains unchanged), but differs by the non-trivial element in the fiber. This adjustment ensures that the classes map to the same element in $\pi_{-1}(B\mathrm{Spin}^{c}(n-1)^{-(L^{\mathrm{Spin}^{c}}_{n-1}\oplus\mathcal{O}_{n-1})})$. This proves sufficiency. ∎
Using this lemma, we observe that the difference
$M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]-[M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]\in\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$
vanishes in $\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$. Consequently, it naturally maps to 0 in $\Omega_{n}^{\mathrm{Spin}}$ (via the map to the base space). Thus, this difference lies in the image of the map from $\mathbb{Z}$} (from the vertical exact sequence). However, since it vanishes in $\pi_{-2}(B\mathrm{Spin}^{c}(n)^{-(L^{\mathrm{Spin}^{c}}_{n}\oplus\mathcal{O}_{n})})$, the integer pre-image must lie in the image of $p$ (from the horizontal structure).
We analyze this situation using the following extended commutative diagram:
[…]
Based on this diagram and previous considerations, we state the reformulation of $M_{1}\sim M_{2}$ depending on the value of $n$:
  • When $n$ is even, $p$ is the zero map. Thus, the map $u$ is injective. Therefore, $M_{1}\sim M_{2}$ is equivalent to
    $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]=[M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]\in\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}}).$
    For even $n$, equality as elements of $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ is equivalent to equality as elements of $\Omega_{n}^{\mathrm{Spin}}$ combined with the equality of their Euler characteristics [9, 6, 8]. Since the condition on the Euler characteristic is clearly satisfied by assumption (both are 0), $M_{1}\sim M_{2}$ is equivalent to $M_{1}]=[M_{2}]\in\Omega_{n}^{\mathrm{Spin}}$.
  • When $n\equiv 1,5,7\pmod{8}$, $p$ is surjective. Therefore, $M_{1}\sim M_{2}$ is equivalent to
    $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]-[M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]\in\mathrm{Im}(t).$
    Due to the exactness of the left vertical column, this is equivalent to $M_{1}]=[M_{2}]\in\Omega_{n}^{\mathrm{Spin}}$.
  • When $n\equiv 3\pmod{8}$, the image of $p$ consists of all even integers. Thus, $M_{1}\sim M_{2}$ is equivalent to
    $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]-[M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]\in t(2{\mathbb{Z}}).$
    Here, since the Euler characteristic of $S^{n+1}$ is 2, the element $2\in{\mathbb{Z}}$ lies in the kernel of $t$ (as the boundary of the disk bundle). Consequently, $t(2{\mathbb{Z}})=0$, and in this case, $M_{1}\sim M_{2}$ is equivalent to
    $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]=[M_{2},{\mathbb{R}}\subset TM_{2}\oplus{\mathbb{R}}]\in\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}}).$
    Furthermore, since $n$ is odd and the Euler characteristic of an $(n+1)$-dimensional closed $\mathrm{Spin}$ manifold cannot be odd, equality in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ is equivalent to equality in $\Omega_{n}^{\mathrm{Spin}}$ combined with the equality of the Kervaire semi-characteristic $\hat{\chi}_{{\mathbb{Z}}/2{\mathbb{Z}}}$ [6, 8].
Summarizing the above:
  • If $n\not\equiv 3\pmod{8}$, then $M_{1}\sim M_{2}$ is equivalent to $M_{1}]=[M_{2}]\in\Omega_{n}^{\mathrm{Spin}}$.
  • If $n\equiv 3\pmod{8}$, then $M_{1}\sim M_{2}$ is equivalent to $M_{1}]=[M_{2}]\in\Omega_{n}^{\mathrm{Spin}}$ and $\hat{\chi}_{{\mathbb{Z}}/2{\mathbb{Z}}}(M_{1})=\hat{\chi}_{{\mathbb{Z}}/2{\mathbb{Z}}}(M_{2})\in{\mathbb{Z}}/2{\mathbb{Z}}$.
This result provides a more precise result than the one presented in [1].

### v2

@@
 Henceforth, we denote the Thom spectrum of a virtual vector bundle $V\to X$ by $X^{V}$.
-Definition 1.1.
+Definition $1.$
 Let $V\to X$ be a virtual vector bundle of rank $n\in{\mathbb{Z}}_{\geq 0}$. We define the monoid denoted by $\Omega(X;V)$, as the set of equivalence classes of triples $(M,g,\Phi)$ consisting of:
@@
   • a continuous map $g:M\to X$, and
-  • a stable vector bundle isomorphism $\Phi:TM\cong g^{*}V$ (i.e., there exists a sufficiently large positive integer $N$ and a bundle isomorphism $\Phi:TM\oplus\mathbb{R}^{N}\cong g^{V}\oplus\mathbb{R}^{N}$).
+  • a stable vector bundle isomorphism $\Phi:TM\cong g^{*}V$ (i.e., there exists a sufficiently large positive integer NN and a bundle isomorphism $\Phi:TM\oplus\mathbb{R}^{N}\cong g^{V}\oplus\mathbb{R}^{N}$).
 The equivalence relation is defined as follows. A direct cobordism between two triples $(M_{0},g_{0},\Phi_{0})$ and $(M_{1},g_{1},\Phi_{1})$ is a triple $(N,h,\Psi)$ consisting of:
@@
 Furthermore, there exists a natural surjective group homomorphism
-$\epsilon:\Omega^{\textnormal{Spin, line}}_{n}\to\Omega^{O}_{n-1}\oplus\Omega^{SO}_{n}$
+$\epsilon:\Omega^{\textnormal{Spin, line}}_{n}\to\Omega^{\mathrm{Pin}^{-}}_{n-1}\oplus\Omega^{\mathrm{Spin}}_{n}$
 defined by $\epsilon([M,\eta])=([s^{-1}(0)],[M])$, where $s$ is a generic section of $\eta$ transverse to the zero-section. The kernel of $\epsilon$ is given by
@@
 As corollaries of our main theorems, we present two applications.
-The first application determines the non-negative integers $n$ for which there exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact Spin manifold $M$ with boundary $S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta$ is orthogonal to the boundary on $S^{n}$.
+The first application determines the non-negative integers $n$ for which there exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact Spin manifold $M$ with boundary $S^{n}$ and a line subbundle $\eta$ of TMTM such that $\eta$ is orthogonal to the boundary on $S^{n}$.
 The second application provides an alternative proof of the main result in [1].
@@
 If the implication holds (i.e., mapping to zero implies being zero), we have $a+c=b+c=0$, which implies $a=a+(b+c)=(a+c)+b=0+b=b$.
-Since $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, there exist a sufficiently large positive integer $N$, a sequence of $n$-dimensional closed manifolds $M=M_{0},M_{1},\ldots,M_{l}$ with $M_{l+1}=\emptyset$, continuous maps $g_{i}:M_{i}\to X$, and bundle isomorphisms $\Phi_{i}$ between $TM_{i}\oplus{\mathbb{R}}^{m+N}$ and $g_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$.
+Since $M,g,\phi]=0\in\Omega(X;\xi-{\mathbb{R}}^{m})$, there exist a sufficiently large positive integer NN, a sequence of $n$-dimensional closed manifolds $M=M_{0},M_{1},\ldots,M_{l}$ with $M_{l+1}=\emptyset$, continuous maps $g_{i}:M_{i}\to X$, and bundle isomorphisms $\Phi_{i}$ between $TM_{i}\oplus{\mathbb{R}}^{m+N}$ and $g_{i}^{*}\xi\oplus{\mathbb{R}}^{N}$.
 Note that we set $g_{0}=g$ and $\Phi_{0}=\phi\oplus{\mathbb{R}}^{N}$.
@@
 Consequently, $2[r\cdot c]=0\in\pi_{n}(GL^{+}(n+1))$ holds.
-Therefore, $a$ can be taken as either 0 or $1$, which completes the proof for this case.
+Therefore, $a$ can be taken as either 0 or 11, which completes the proof for this case.
 The remaining case is when $n$ is odd and $a$ is negative. First, similar to the positive case, we can reduce this to the case a=1a=-1. We begin by proving the following lemma.
-Lemma 3.3.
+Lemma $3.$
 For a continuous map $f:S^{n}\to GL(n+1)$, let $\chi_{f}$ denote the degree of the map $S^{n}\to S^{n}$ obtained by composing $f$ with the natural map $GL(n+1)\to S^{n}$. If $n$ is odd, then
@@
 The strategy of this proof is analogous to the first half of the proof of Theorem 4.4 in [3].
-Suppose $M_{0}\sim M_{1}$ via a cobordism $N$ between them.
-We consider constructing a vector field on $N$ that points outwards along the boundaries $M_{0}$ and $M_{1}$.
-If such a vector field is nowhere zero, we can compose the trivialization of $TN$ with the reflection defined by this vector field.
+Suppose $M_{0}\sim M_{1}$ via a cobordism NN between them.
+We consider constructing a vector field on NN that points outwards along the boundaries $M_{0}$ and $M_{1}$.
+If such a vector field is nowhere zero, we can compose the trivialization of TNTN with the reflection defined by this vector field.
 This allows us to swap the inward and outward directions of the $\oplus{\mathbb{R}}$ factor at both $M_{0}$ and $M_{1}$, thereby establishing symmetry.
-If the Euler characteristic of $N$ is non-zero, such a vector field may not exist globally; however, one can always construct such a vector field away from a finite set of points.
-By excising open neighborhoods of these points from $N$, we obtain the relation $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k}S^{n}$.
+If the Euler characteristic of NN is non-zero, such a vector field may not exist globally; however, one can always construct such a vector field away from a finite set of points.
+By excising open neighborhoods of these points from NN, we obtain the relation $M_{1}\sim M_{0}\cup\bigcup_{j=1}^{k}S^{n}$.
 It remains to show that these $S^{n}$ components are null-cobordant without switching the inward/outward direction of $\oplus{\mathbb{R}}$.
@@
 Let $X^{n+1}$ denote $S^{\frac{n+1}{2}}\times S^{\frac{n+1}{2}}$ if $n+1$ is divisible by 4, and $S^{\frac{n-1}{2}}\times S^{\frac{n+3}{2}}$ if n1n-1 is divisible by 4.
-Note that $X^{n+1}$ is a stably trivial, connected, closed $(n+1)$-dimensional manifold with Euler characteristic $4$ (this proof requires $n\neq 1$, as connectedness fails otherwise).
+Note that $X^{n+1}$ is a stably trivial, connected, closed $(n+1)$-dimensional manifold with Euler characteristic 44 (this proof requires $n\neq 1$, as connectedness fails otherwise).
 Since a stably trivial closed manifold is almost parallelizable, $TX^{n+1}$ admits a trivialization on $X^{n}$ minus a single point.
@@
 Since $X^{n+1}$ is stably trivial, this element maps to 0 under the stabilization map $S^{n},GL(n+1)]\to[S^{n},GL(n+2)]$.
-Applying the Poincaré-Hopf theorem to $X^{n+1}$, the image of this element under the map $S^{n},GL(n+1)]\to[S^{n},S^{n}]$ coincides with the Euler characteristic of $X^{n+1}$, which is $4$.
-By the lemma stated earlier, swapping the outward and inward directions of $\oplus{\mathbb{R}}$ transforms an element of degree $4$ into an element of degree 24=22-4=-2 in $S^{n},GL(n+1)]$.
-Consequently, an $S^{n}$ carrying an element of degree 2-2 (corresponding to $r\cdot c]$ since the degree of $r\cdot c]$ is $2$) is bounded by $X^{n+1}\setminus\mathrm{int}(D^{n+1})$.
+Applying the Poincaré-Hopf theorem to $X^{n+1}$, the image of this element under the map $S^{n},GL(n+1)]\to[S^{n},S^{n}]$ coincides with the Euler characteristic of $X^{n+1}$, which is 44.
+By the lemma stated earlier, swapping the outward and inward directions of $\oplus{\mathbb{R}}$ transforms an element of degree 44 into an element of degree 24=22-4=-2 in $S^{n},GL(n+1)]$.
+Consequently, an $S^{n}$ carrying an element of degree 2-2 (corresponding to $r\cdot c]$ since the degree of $r\cdot c]$ is 22) is bounded by $X^{n+1}\setminus\mathrm{int}(D^{n+1})$.
 In conclusion, all such $S^{n}$ are null-cobordant without requiring a change in the boundary normal direction.
@@
 Since the cobordism group $\Omega^{\prime}(BSO(2);L^{SO}_{2}\oplus{\mathbb{R}}^{2},1)$ is trivial, every element is null-cobordant (i.e., it bounds some manifold).
-In particular, given a pair $(M,X)$ consisting of a closed oriented $3$-manifold $M$ and a nowhere-vanishing vector field $X$ on $M$, there exists a triple $(N,Y_{1},Y_{2})$ consisting of a compact oriented $4$-manifold NN and two linearly independent vector fields $Y_{1},Y_{2}$ on $N$ such that:
+In particular, given a pair $(M,X)$ consisting of a closed oriented 33-manifold $M$ and a nowhere-vanishing vector field $X$ on $M$, there exists a triple $(N,Y_{1},Y_{2})$ consisting of a compact oriented 44-manifold NN and two linearly independent vector fields $Y_{1},Y_{2}$ on NN such that:
   • $\partial N\cong M$;
   • $Y_{1}|_{M}=X$, and $Y_{2}|_{M}$ is the outward-pointing normal vector field.
-This result is stronger than the fact that the $3$-dimensional oriented cobordism group is trivial; it means that we can impose additional structure on the bounding manifolds.
+This result is stronger than the fact that the 33-dimensional oriented cobordism group is trivial; it means that we can impose additional structure on the bounding manifolds.
 4. The proof of Theorem B
@@
 Then $\pi_{-m}(BO(n)^{-L_{n}^{O}})$ is isomorphic to the (nm)(n-m)-dimensional cobordism group, $\Omega_{n-m}^{O}$.
-A similar theorem holds for $BSO$, $B\mathrm{Spin}$, and $B\mathrm{Pin}^{-}$.
+A similar theorem holds for BSOBSO, $B\mathrm{Spin}$, and $B\mathrm{Pin}^{-}$.
 Here, for convenience, we define $\Omega_{n-m}^{O}=0$ when $n<m$.
@@
 Since these correspondences are inverses of each other, the first part of Theorem C2 is established.
-The second part of the proof follows the same logic as in Theorem C1, with $SO$ replaced by $\mathrm{Spin}$ and $O$ replaced by $\mathrm{Pin}^{-}$, up to the final step.
+The second part of the proof follows the same logic as in Theorem C1, with SOSO replaced by $\mathrm{Spin}$ and $O$ replaced by $\mathrm{Pin}^{-}$, up to the final step.
 It remains to determine the values of $n$ for which the map $p:\pi_{-1}(B\mathrm{Pin}^{-}(n+1)^{-(L^{\mathrm{Pin}^{-}}_{n+1}\oplus\mathrm{det}L^{\mathrm{Pin}^{-}}_{n+1})})\to\Omega^{\mathrm{Pin}^{-}}_{n}\oplus\Omega^{\mathrm{Spin}}_{n+1}\to{\mathbb{Z}}/2{\mathbb{Z}}$ is the zero map.
@@
 Since $2e=0$ in odd dimensions (for the vector bundle, i.e., when $n+1$ is odd), $p$ is the zero map when $n$ is even.
-When $n$ is odd, considering the element $S^{n+1},{\mathbb{R}}^{2}\subset TS^{n+1}\oplus{\mathbb{R}}^{2}]$, its image under $p$ is the Euler characteristic of $S^{n+1}$, which is $2$.
-Thus, $2$ is contained in the image of $p$.
-It remains only to consider the parity of the values of $p$.
+When $n$ is odd, considering the element $S^{n+1},{\mathbb{R}}^{2}\subset TS^{n+1}\oplus{\mathbb{R}}^{2}]$, its image under $p$ is the Euler characteristic of $S^{n+1}$, which is 22.
+Thus, 22 is contained in the image of pp.
+It remains only to consider the parity of the values of pp.
 The mod 2 reduction of the Euler class $e$ is $w_{n+1}$, and we have
@@
 Theorem 6.1.
-Consider the following condition for a non-negative integer $n$: ”There exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact $\mathrm{Spin}$ manifold $M$ with boundary $\partial M=S^{n}$ and a line subbundle $\eta$ of $TM$ such that $\eta|_{S^{n}}=(TS^{n})^{\perp}$ (i.e., $\eta$ is orthogonal to the boundary along $S^{n}$).”
+Consider the following condition for a non-negative integer $n$: ”There exists a pair $(M,\eta)$ consisting of an $(n+1)$-dimensional compact $\mathrm{Spin}$ manifold $M$ with boundary $\partial M=S^{n}$ and a line subbundle $\eta$ of TMTM such that $\eta|_{S^{n}}=(TS^{n})^{\perp}$ (i.e., $\eta$ is orthogonal to the boundary along $S^{n}$).”
 Then, this condition holds if and only if $n\equiv 0,2,6,7\pmod{8}$.
@@
 When $n$ is odd, according to [4], the given condition is equivalent to the Euler characteristic of $M$ being zero.
-Consequently, the proposition is equivalent to the existence of a closed $(n+1)$-dimensional $\mathrm{Spin}$ manifold with Euler characteristic $1$.
+Consequently, the proposition is equivalent to the existence of a closed $(n+1)$-dimensional $\mathrm{Spin}$ manifold with Euler characteristic 11.
 As mentioned previously, such manifolds exist only when $n\equiv 7\pmod{8}$ (i.e., when the dimension $n+1$ is divisible by 8), in which case $M=\mathbb{HP}^{\frac{n+1}{4}}\setminus\mathrm{int}(D^{n+1})$ satisfies the condition.
@@
   • $W$ is a compact $(n+1)$-dimensional $\mathrm{Spin}$ manifold with boundary $\partial W=M_{1}\cup M_{2}$.
-  • $\xi$ is a rank 2 oriented subbundle of $TW$.
+  • $\xi$ is a rank 2 oriented subbundle of TWTW.
   • Restricted to $M_{i}$, $\xi$ decomposes as the direct sum of the trivial line bundle $\mathbb{R}$} (which is orthogonal to the boundary) and a trivial line subbundle $\eta$ of $TM_{i}$.
@@
 If this element is 0, then $M_{1}\sim M_{2}$ holds by the definition of the equivalence relation.
-If this element is the image of the non-zero element of ${\mathbb{Z}}/2{\mathbb{Z}}$, we can adjust it using the image of $1$ under the map ${\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$.
+If this element is the image of the non-zero element of ${\mathbb{Z}}/2{\mathbb{Z}}$, we can adjust it using the image of 11 under the map ${\mathbb{Z}}/2{\mathbb{Z}}\to\pi_{1}(B\mathrm{Spin}(n-1)^{-L^{\mathrm{Spin}}_{n-1}})$.
 Specifically, we can replace $M_{1},\eta_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ with another element $M_{1},\eta^{\prime}_{1}\oplus{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ such that it maps to the same element $M_{1},{\mathbb{R}}\subset TM_{1}\oplus{\mathbb{R}}]$ in $\pi_{0}(B\mathrm{Spin}(n)^{-L^{\mathrm{Spin}}_{n}})$ (the base class remains unchanged), but differs by the non-trivial element in the fiber.